# Adjusting Model Size in Continual Gaussian Processes: How Big is Big Enough?

### v4

@@ B.2 Conditional Variance
The implementation of the Conditional Variance method is presented in Algorithm 5. This method uses the “greedy variance” strategy that iteratively chooses the location of the next inducing point. As a stopping criterion, it uses the quantity $\operatorname{tr}\left(\mathbf{K_{ff}}-\mathbf{Q_{ff}}\right)$. In this algorithm, new inducing points are no longer added once $\operatorname{tr}\left(\mathbf{K_{ff}}-\mathbf{Q_{ff}}\right)$ falls below a chosen tolerance value $\eta$. Although, this approach was mentioned in Burt et al. (2020), this stopping criterion has not yet been tested in the literature. The hyperparameter $\eta$ is determined by the user.
@@ B.3 Online Inducing Point Selection (OIPS)
Galy-Fajou & Opper (2021) introduced the Online Inducing Points Selection (OIPS) algorithm, which iteratively adds points from $\mathbf{X}_{n}$ to the set of inducing points. The algorithm assesses the impact of each new point on the existing inducing set, based on a covariance threshold. A point $\mathbf{x}$ is added if the maximum value of $\mathbf{k}_{\mathbf{u}}(\mathbf{x})$ falls below a user-defined threshold $\rho$. While the original algorithm implicitly assumes unit kernel variance, Algorithm 6 presents our implementation where we scale the threshold $\rho$ by the kernel variance $\sigma^{2}_{f}$ to $\bar{\rho}=\sigma^{2}_{f}\rho$. This modification ensures the threshold comparison remains valid for arbitrary kernel variances.
@@
  Input: $\mathbf{X}_{n}=\{\mathbf{x}_{i}\}_{i=1}^{N_{n}}$, $\mathbf{Z}_{o}=\{\mathbf{z}_{m}\}_{m=1}^{M_{\mathbf{a}}}$, kernel function $k(\cdot,\cdot|\theta_{n})$, kernel hyperparameters $\theta_{n}$ (including variance $\sigma_{f}^{2}$), acceptance threshold $10<\rho<$ 1.
@@ Appendix C Proof of Guarantee
where $\Psi=\int q_{n}(f)\log\frac{q_{n}^{*}(f)}{p(f|\mathbf{y}_{n},\mathbf{y}_{o})}\mathrm{d}f$ and $q_{n}^{*}(f)=\mathcal{Z}^{-1}q_{o}(f)p(\mathbf{y}_{n}|f)$ represents the variational distribution associated with the optimal lower bound =^$\mathcal{L}^{*}=\widehat{\mathcal{L}}(N_{n}+M_{\mathbf{a}})$, with $\mathcal{Z}$ denoting the marginal likelihood that normalises $q_{n}^{*}(f)$.
@@
We cease the addition of points when $\widehat{\mathcal{U}}(M)-\widehat{\mathcal{L}}(M_{\mathbf{b}})<\alpha$. Since $\widehat{\mathcal{U}}(M)\geq\mathcal{L}^{*}$, then ^$-\widehat{\mathcal{L}}(M_{\mathbf{b}})<\alpha-\widehat{\mathcal{U}}(M)<\alpha-\mathcal{L}^{*}$. Eq.(4) can be bounded as:
@@
where $\Phi=-\int q_{n}(f)\log\frac{p(f|\mathbf{y}_{o},\theta_{o})}{q_{o}(f)}\text{d}f$. Let $q_{n}^{*}(f)=\mathcal{Z}^{-1}q_{o}(f)p(\mathbf{y}_{n}|f)$ be the variational distribution associated with =^$\mathcal{L}^{*}=\widehat{\mathcal{L}}(N_{n}+M_{\bf a})$. Then, by expanding the true posterior and multiplying by the variational distributions $q_{n}^{*}(f)$ on both sides of the fraction inside the log, we obtain:
@@
[…]
@@
$\operatorname{KL}[q_{n}(f)\mid\mid p(f|\mathbf{y}_{o},\mathbf{y}_{n},\theta_{o})]$ $\leq\int q_{n}(f)\log\frac{q_{n}^{*}(f)}{p(f|\mathbf{y}_{n},\mathbf{y}_{o},\theta_{n})}\frac{q_{n}(f)}{q_{n}(f)}\text{d}f+\alpha$ (19)
$\operatorname{KL}[q_{n}(f)\mid\mid p(f|\mathbf{y}_{o},\mathbf{y}_{n},\theta_{n})]$ $\leq\operatorname{KL}[q_{n}(f)\mid\mid p(f|\mathbf{y}_{o},\mathbf{y}_{n},\theta_{n})]-\operatorname{KL}[q_{n}(f)\mid\mid q_{n}^{*}(f)]+\alpha$
@@ D.1 Model size and data distribution
We use $N=500$ observations uniformly distributed from 00 to $10$. The data is ordered and divided into ten batches.
@@
We simulate a scenario where small batches of data are received but the data is distributed across the input space. We use $N=150$ observations uniformly distributed from 00 to $10$. The data is shuffled and divided into ten batches.
@@ D.3 UCI datasets
These experiments were performed on an Nvidia RTX 6000’s GPU on a high-performance computing cluster. We used a Squared Exponential kernel with hyperparameters initialised to 1 for all datasets. The noise variance was initialised to 0.1. We consider six UCI (Dua & Graff, 2017) datasets of different characteristics: Concrete (1030, 8), Skillcraft (3338, 19), Kin8nm (8192, 8), Naval (11934, 14), Elevators (16599, 18), and Bike (17379, 17). The smaller datasets $(<12000)$ were divided into 20 batches, and the larger $(>12000)$ into 50 batches. The batches were created by sorting the data points by the first dimension.
@@ D.4 Magnetic anomalies
The data used in this experiment is obtained from Solin et al. (2018) and is available on GitHub. The objective of this task is to detect local anomalies in the Earth’s magnetic field online, caused by the presence of bedrock and magnetic materials in indoor building structures. For this purpose, a small robot with a 3-axis magnetometer moves around an indoor space of approximately 6 meters by 6 meters and measures the magnetic field strength. Out of the 9 available trajectories, we use trajectories 1, 2, 3, 4, and 5 (with $n=8875,9105,9404,7332,8313$, respectively) for the experiments. Specifically, we use trajectories 1, 2, 4 and 5 for the first experiment and trajectory 3 for the second.
We use the experimental setup proposed in Chang et al. (2023a). The proposed model applies a GP prior to magnetic field strength, given by $\mathcal{G}\mathcal{P}\left(0,\sigma_{0}^{2}+\kappa_{\sigma^{2},\ell}^{\mathrm{Mat}}\left(\mathbf{x},\mathbf{x}^{\prime}\right)\right)$ (in $\mu\mathrm{T}$ ), where the kernel consist of a constant kernel and a Matérn-$\nu/2$ kernel. The model assumes the spatial domain is affected by Gaussian noise with a variance $\sigma_{\mathrm{n}}^{2}$. The initial variance for the constant kernel is set to $500$, and the Gaussian likelihood is initialised with a noise variance of $0.1$.
@@ D.5 Model size and Data Distribution with Larger Data
We reproduce the three continual learning scenarios from App. D.1, increasing the dataset size to 50 0005000050\,00050 000 data points and the input dimensionality to three, and compare how the three methods (VIPS, OIPS and Conditional Variance) scale in this setting. This setup aims to simulate practical scenarios where an exact GP would be infeasible, and a sequential approach is required instead. We use a Squared Exponential kernel initialised with lengthscale 1.0 and variance 1.0. The noise variance is set to 0.2. We evaluate all methods at the operating points corresponding to the 10% RMSE thresholds determined from the UCI experiments (App. D.3). The data is divided into $50$ batches of $1000$ points each.
@@
Data points are sampled uniformly from $[0,20]^{3}$, with batches ordered such that new regions of the input space are discovered gradually. Model size is expected to grow linearly as more data is observed.
@@
Data points are again sampled uniformly from $[0,20]^{3}$, but batches are shuffled at random. This represents a setting where the entire input space is observed from the start. As a result, the novelty of each batch decreases over time, and the model size is expected to plateau.
@@
Most data is concentrated in a narrow region, with occasional outliers. Ninety per cent of the points are drawn from a truncated Cauchy distribution on $[5,15]^{3}$ centred at 10, and the remainder uniformly from $[0,40]^{3}$. This mimics a scenario where data is dense but includes some outliers. Therefore, the model size should remain small until the region with the outliers is observed.
@@ Appendix E Derivation the optimal form of variational distribution
The derivations follow Bui et al. (2017). We start by rewriting the online lower bound,
@@
$q_{\mathrm{opt}}(\mathbf{b})\propto p(\mathbf{b}|\theta_{n})\mathcal{N}\left(\hat{\mathbf{y}},\mathbf{K}_{\hat{\mathbf{f}}\mathbf{b}}\mathbf{K}_{\mathbf{b}\mathbf{b}}^{-1}\mathbf{b},\Sigma_{\hat{\mathbf{y}}}\right)$ (23)
@@
^$\widehat{\mathcal{L}}=\log\mathcal{N}\left(\hat{\mathbf{y}};\mathbf{0},\mathbf{K}_{\hat{\mathbf{f}}\mathbf{b}}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{K}_{\mathbf{b}\hat{\mathbf{f}}}+\Sigma_{\hat{\mathbf{y}}}\right)+\Delta_{\mathbf{a}}-\frac{1}{2}\operatorname{tr}\left[\mathbf{D}_{\mathbf{a}}^{-1}(\mathbf{K}_{\mathbf{aa}}-\mathbf{Q}_{\mathbf{aa}})\right]-\frac{1}{2\sigma^{2}}\operatorname{tr}(\mathbf{K}_{\mathbf{ff}}-\mathbf{Q}_{\mathbf{ff}}).$ (26)
@@ E.1 Derivation of $\mathcal{L}^{*}$
[…]
@@
$\operatorname{tr}\left(\Sigma_{\mathbf{\hat{y}}}^{-1}\left(\mathbf{K}_{\mathbf{\hat{f}\hat{f}}}-\mathbf{K}_{\mathbf{\hat{f}b}}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{K}_{\mathbf{b\hat{f}}}\right)\right)$ (29)
@@
which recovers the expression for ^$\widehat{\mathcal{L}}-\Delta_{\mathbf{a}}$
@@ Appendix F Online Upper Bound Implementation
[…]
@@ F.1 Determinant term
$\log|\mathbf{K}_{\mathbf{\hat{f}b}}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{K}_{\mathbf{b\hat{f}}}+\Sigma_{\mathbf{\hat{y}}}|=$ $\log|\Sigma_{\mathbf{\hat{y}}}|+\log|\mathbf{I}+\mathbf{L}_{\mathbf{b}}^{-1}\mathbf{K}_{\mathbf{b\hat{f}}}\Sigma_{\mathbf{\hat{y}}}^{-1}\mathbf{K}_{\mathbf{\hat{f}b}}\mathbf{L}_{\mathbf{b}}^{-T}|$ (31)
@@ F.2 Quadratic term
$-\frac{1}{2}\hat{\mathbf{y}}^{T}\left(\mathbf{K}_{\mathbf{\hat{f}b}}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{K}_{\mathbf{b\hat{f}}}+t\mathbf{I}+\Sigma_{\mathbf{\hat{y}}}\right)^{-1}\hat{\mathbf{y}}.$

### v5

@@ B.2 Conditional Variance
The implementation of the Conditional Variance method is presented in Algorithm 5. This method uses the “greedy variance” strategy that iteratively chooses the location of the next inducing point. As a stopping criterion, it uses the quantity $\operatorname{tr}\left(\mathbf{K_{ff}}-\mathbf{Q_{ff}}\right)$. In this algorithm, new inducing points are no longer added once $\operatorname{tr}\left(\mathbf{K_{ff}}-\mathbf{Q_{ff}}\right)$ falls below a chosen tolerance value $\eta$. Although, this approach was mentioned in Burt2020, this stopping criterion has not yet been tested in the literature. The hyperparameter $\eta$ is determined by the user.
@@ B.3 Online Inducing Point Selection (OIPS)
Galy-Fajou2021-tu introduced the Online Inducing Points Selection (OIPS) algorithm, which iteratively adds points from $\mathbf{X}_{n}$ to the set of inducing points. The algorithm assesses the impact of each new point on the existing inducing set, based on a covariance threshold. A point $\mathbf{x}$ is added if the maximum value of $\mathbf{k}_{\mathbf{u}}(\mathbf{x})$ falls below a user-defined threshold $\rho$. While the original algorithm implicitly assumes unit kernel variance, Algorithm 6 presents our implementation where we scale the threshold $\rho$ by the kernel variance $\sigma^{2}_{f}$ to $\bar{\rho}=\sigma^{2}_{f}\rho$. This modification ensures the threshold comparison remains valid for arbitrary kernel variances.
@@
Input: $\mathbf{X}_{n}=\{\mathbf{x}_{i}\}_{i=1}^{N_{n}}$, $\mathbf{Z}_{o}=\{\mathbf{z}_{m}\}_{m=1}^{M_{\mathbf{a}}}$, kernel function $k(\cdot,\cdot|\theta_{n})$, kernel hyperparameters $\theta_{n}$ (including variance $\sigma_{f}^{2}$), acceptance threshold $0<\rho<1$.
@@ Appendix C Proof of Guarantee
where $\Psi=\int q_{n}(f)\log\frac{q_{n}^{*}(f)}{p(f|\mathbf{y}_{n},\mathbf{y}_{o})}\mathrm{d}f$ and $q_{n}^{*}(f)=\mathcal{Z}^{-1}q_{o}(f)p(\mathbf{y}_{n}|f)$ represents the variational distribution associated with the optimal lower bound $\mathcal{L}^{*}=\widehat{\mathcal{L}}(N_{n}+M_{\mathbf{a}})$, with $\mathcal{Z}$ denoting the marginal likelihood that normalises $q_{n}^{*}(f)$.
@@
We cease the addition of points when $\widehat{\mathcal{U}}(M)-\widehat{\mathcal{L}}(M_{\mathbf{b}})<\alpha$. Since $\widehat{\mathcal{U}}(M)\geq\mathcal{L}^{*}$, then $\widehat{\mathcal{L}}(M_{\mathbf{b}})<\alpha-\widehat{\mathcal{U}}(M)<\alpha-\mathcal{L}^{*}$. Eq.(4) can be bounded as:
@@
where $\Phi=-\int q_{n}(f)\log\frac{p(f|\mathbf{y}_{o},\theta_{o})}{q_{o}(f)}\text{d}f$. Let $q_{n}^{*}(f)=\mathcal{Z}^{-1}q_{o}(f)p(\mathbf{y}_{n}|f)$ be the variational distribution associated with $\mathcal{L}^{*}=\widehat{\mathcal{L}}(N_{n}+M_{\bf a})$. Then, by expanding the true posterior and multiplying by the variational distributions $q_{n}^{*}(f)$ on both sides of the fraction inside the log, we obtain:
@@
$=\int q_{n}(f)\log\frac{\cancel{q_{o}(f)}p(\mathbf{y}_{o}|\theta_{o})}{p(\mathbf{y}_{o}|f)p(f|\theta_{o})}\frac{q_{n}^{*}(f)}{\mathcal{Z}^{-1}\cancel{q_{o}(f)}p(\mathbf{y}_{n}|f)}\text{d}f$
@@
$\operatorname{KL}[q_{n}(f)\mid\mid p(f|\mathbf{y}_{o},\mathbf{y}_{n},\theta_{n})]$ $\leq\int q_{n}(f)\log\frac{q_{n}^{*}(f)}{p(f|\mathbf{y}_{n},\mathbf{y}_{o},\theta_{n})}\frac{q_{n}(f)}{q_{n}(f)}\text{d}f+\alpha$ (19)
$\operatorname{KL}[q_{n}(f)\mid\mid p(f|\mathbf{y}_{o},\mathbf{y}_{n},\theta_{n})]$ $\leq\operatorname{KL}[q_{n}(f)\mid\mid p(f|\mathbf{y}_{n},\mathbf{y}_{o},\theta_{n})]-\operatorname{KL}[q_{n}(f)\mid\mid q_{n}^{*}(f)]+\alpha$
@@ D.1 Model size and data distribution
We use $N=500$ observations uniformly distributed from 0 to $10$. The data is ordered and divided into ten batches.
@@
We simulate a scenario where small batches of data are received but the data is distributed across the input space. We use $N=150$ observations uniformly distributed from 0 to $10$. The data is shuffled and divided into ten batches.
@@ D.3 UCI datasets
These experiments were performed on an Nvidia RTX 6000’s GPU on a high-performance computing cluster. We used a Squared Exponential kernel with hyperparameters initialised to 1 for all datasets. The noise variance was initialised to 0.1. We consider six UCI (UCI) datasets of different characteristics: Concrete (1030, 8), Skillcraft (3338, 19), Kin8nm (8192, 8), Naval (11934, 14), Elevators (16599, 18), and Bike (17379, 17). The smaller datasets $(<12000)$ were divided into 20 batches, and the larger $(>12000)$ into 50 batches. The batches were created by sorting the data points by the first dimension.
@@ D.4 Magnetic anomalies
The data used in this experiment is obtained from Solin2018 and is available on GitHub. The objective of this task is to detect local anomalies in the Earth’s magnetic field online, caused by the presence of bedrock and magnetic materials in indoor building structures. For this purpose, a small robot with a 3-axis magnetometer moves around an indoor space of approximately 6 meters by 6 meters and measures the magnetic field strength. Out of the 9 available trajectories, we use trajectories 1, 2, 3, 4, and 5 (with $n=8875,9105,9404,7332,8313$, respectively) for the experiments. Specifically, we use trajectories 1, 2, 4 and 5 for the first experiment and trajectory 3 for the second.
We use the experimental setup proposed in Chang2023. The proposed model applies a GP prior to magnetic field strength, given by $\mathcal{G}\mathcal{P}\left(0,\sigma_{0}^{2}+\kappa_{\sigma^{2},\ell}^{\mathrm{Mat}}\left(\mathbf{x},\mathbf{x}^{\prime}\right)\right)$ (in $\mu\mathrm{T}$ ), where the kernel consist of a constant kernel and a Matérn-$\nu/2$ kernel. The model assumes the spatial domain is affected by Gaussian noise with a variance $\sigma_{\mathrm{n}}^{2}$. The initial variance for the constant kernel is set to $500$, and the Gaussian likelihood is initialised with a noise variance of $0.1$.
@@ D.5 Model size and Data Distribution with Larger Data
We reproduce the three continual learning scenarios from App. D.1, increasing the dataset size to 50 00050\,000 data points and the input dimensionality to three, and compare how the three methods (VIPS, OIPS and Conditional Variance) scale in this setting. This setup aims to simulate practical scenarios where an exact GP would be infeasible, and a sequential approach is required instead. We use a Squared Exponential kernel initialised with lengthscale 1.0 and variance 1.0. The noise variance is set to 0.2. We evaluate all methods at the operating points corresponding to the 10% RMSE thresholds determined from the UCI experiments (App. D.3). The data is divided into $50$ batches of $1000$ points each.
@@
Data points are sampled uniformly from $0,20]^{3}$, with batches ordered such that new regions of the input space are discovered gradually. Model size is expected to grow linearly as more data is observed.
@@
Data points are again sampled uniformly from $0,20]^{3}$, but batches are shuffled at random. This represents a setting where the entire input space is observed from the start. As a result, the novelty of each batch decreases over time, and the model size is expected to plateau.
@@
Most data is concentrated in a narrow region, with occasional outliers. Ninety per cent of the points are drawn from a truncated Cauchy distribution on $5,15]^{3}$ centred at 10, and the remainder uniformly from $0,40]^{3}$. This mimics a scenario where data is dense but includes some outliers. Therefore, the model size should remain small until the region with the outliers is observed.
@@ Appendix E Derivation the optimal form of variational distribution
The derivations follow Bui2017. We start by rewriting the online lower bound,
@@
[…]
@@
$\widehat{\mathcal{L}}=\log\mathcal{N}\left(\hat{\mathbf{y}};\mathbf{0},\mathbf{K}_{\hat{\mathbf{f}}\mathbf{b}}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{K}_{\mathbf{b}\hat{\mathbf{f}}}+\Sigma_{\hat{\mathbf{y}}}\right)+\Delta_{\mathbf{a}}-\frac{1}{2}\operatorname{tr}\left[\mathbf{D}_{\mathbf{a}}^{-1}(\mathbf{K}_{\mathbf{aa}}-\mathbf{Q}_{\mathbf{aa}})\right]-\frac{1}{2\sigma^{2}}\operatorname{tr}(\mathbf{K}_{\mathbf{ff}}-\mathbf{Q}_{\mathbf{ff}}).$ (26)
@@ E.1 Derivation of $\mathcal{L}^{*}$
$\mathcal{L}^{*}=\log\mathcal{N}\left(\hat{\mathbf{y}};\,\mathbf{0}\,,\mathbf{K}_{\hat{\mathbf{f}}\hat{\mathbf{f}}}+\Sigma_{\hat{y}}\right)+\Delta_{\mathbf{a}},$ (27)
@@
[…]
@@
which recovers the expression for $\widehat{\mathcal{L}}-\Delta_{\mathbf{a}}$ .
@@ Appendix F Online Upper Bound Implementation
This term is an upper bound for the first term of $\mathcal{L}^{*}=\log\mathcal{N}\left(\hat{\mathbf{y}};\,\mathbf{0}\,,\mathbf{K}_{\hat{\mathbf{f}}\hat{\mathbf{f}}}+\Sigma_{\hat{y}}\right)+\Delta_{\mathbf{a}}$.
@@ F.1 Determinant term
[…]
@@ F.2 Quadratic term
$\frac{1}{2}\hat{\mathbf{y}}^{T}\left(\mathbf{K}_{\mathbf{\hat{f}b}}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{K}_{\mathbf{b\hat{f}}}+t\mathbf{I}+\Sigma_{\mathbf{\hat{y}}}\right)^{-1}\hat{\mathbf{y}}$.